I need this solved using a proportion
A truck can pump 800 gallons of oil in 35 minutes. How much oil can it pump in 2 hours?
: I need this solved using a proportion
A truck can pump 800 gallons of oil in 35 minutes. How much oil can it pump in 2 hours?

Answer by nerdybill(1129)

A truck can pump 800 gallons of oil in 35 minutes. How much oil can it pump in 2 hours?
.
From:"A truck can pump 800 gallons of oil in 35 minutes." we get the proportion:
800/35
.
Since the above proportion is in "minutes" we must convert 2 hours into minutes:
2 hours = 120 mins
.
Therefore, your final proportion to solve is:
If x = oil pumped in 2 hours
then
800/35 = x/120
(800/35)120 = x
2742.857 gallons = x